OpenStudy (anonymous):

A circle is centered at (7, 8) and has a radius of 11. Which of the following is the equation for this circle?

OpenStudy (anonymous):

(x − 7)2 + (y − 8)2 = 121 (x − 7)2 + (y − 8)2 = 11 (x + 7)2 + (y + 8)2 = 121 (x + 7)2 + (y + 8)2 = 11

OpenStudy (anonymous):

equation for a circle: (x-h)^2 + (y-k)^2 = r^2 center = 7,8 = h,k (x - 7)^2 + (y - 8)^2 = 121

OpenStudy (anonymous):

omg thank you so much

OpenStudy (anonymous):

thanks